the gauge pressure in your tires is 30psi when the temperature is 35c. What is the gauge pressure when the temperature becomes -40c?
Answer #1

using the equation

P1/T1 = P2/T2

30/(273 +35) = P2/(273 - 40)

P2 = 22.7 psi

so the gauge pressure at -40 C is 22.7 psi